OpenStudy (anonymous):

Graph the line that passes through the point (6, -8) and has slope -1/3.

OpenStudy (anonymous):

y+8=-1/3(x -6)

OpenStudy (anonymous):

rearanging: y =-1/3x -6

OpenStudy (kymber):

First you want to find B, by plugging the point into y=mx+b equation and solving. B is the y-intercept, the starting point on the. Once you've found that, you can graph the line.